Consider a wire of area A = 1m² initially far from any sources of magnetic field. A strong bar magnet with B = 1T near its poles is suddenly inserted into the loop, filling the loop completely. If the time required to insert the magnet is t = 0.1s , what is the approximate value of the induced emf in the loop?


1
Expert's answer
2021-11-24T12:00:55-0500
"\Epsilon=\dfrac{\Delta\Phi_B}{\Delta t}=\dfrac{BA}{t}=\dfrac{1\ T\times1\ m^2}{0.1\ s}=10\ V."